Hello I am being asked a question that I do not understand, It asks to Solve a system of equations to find the two numbers described.
The sum of two numbers is 17 and their difference is 29

Answer by checkley71(8403)

X+Y=17 OR X=17-Y NOW SUBSTITUTE (17-Y) FOR X IN THE FOLLOWING EQUATION.
X-Y=29
(17-Y)-Y=29
-2Y=29-17
-2Y=12
Y=12/-2
Y=-6 ANSWER.
X-6=17
X=17+6
X=23 ANSWER.
23-(-6)=29
23+6=29
29=29